# Chapter 1      Introduction to IoT

### 09 Dec 2023 v1

@@ 1.6.4 Smart Home (SH)
The adoption of smart home technology has seen a significant and continuous increase in recent years. As technologies such as sensors, actuators, and wireless networks have matured, people’s trust in these advancements has grown, leading to a greater reliance on technology to enhance their quality of life.
In a smart home, various types of sensors are strategically deployed, each with its specific function. Smart homes simplify daily tasks for users, proving especially beneficial for those prone to forgetting routine actions like locking doors or turning off appliances. From smart door locks to the maintenance of household items like coffee machines, heaters, and smart bulbs, and even the use of surveillance cameras for enhanced security, smart homes offer a wide range of possibilities. Furthermore, users can control these devices through voice commands and remotely monitor their home equipment. Smart homes contribute to improved energy efficiency by automatically turning off devices not in use and notifying users of any unusual incidents.
MavHome, short for “Managing An Intelligent Versatile Home” [5], represents a smart home system that functions as an intelligent agent. It employs prediction algorithms to perform various tasks in response to user-initiated events. As for energy conservation, an intelligent home achieves energy conservation through the utilization of sensors and the context-aware capabilities of IoT. Data gathered by these sensors are transmitted to a context aggregator, which then forwards the data to a context-aware service engine. This engine analyzes the data and determines appropriate actions. For instance, it may decide to turn off the air conditioning if the temperature is too cold, shut off the gas supply in case of a detected leak, or switch off the lights when there are no occupants at home [124].
@@ 1.6.5 Smart Manufacturing System (SMS)
With the development and evolution of IoT (Internet of Things), Industrial IoT, Artificial Intelligence, and Cyber-Physical Systems, many countries have opted to transform their manufacturing systems into smart manufacturing systems. Through the integration of smart technologies, these systems facilitate a rapid and extensive flow of data within and among manufacturing processes. Equipped with these data and employing advanced information and communication technology, smart manufacturing systems possess the capability to swiftly respond to global demands, efficiently utilise materials, energy, and labour resources, and deliver customised products on time [125].
What sets the smart manufacturing model apart from other manufacturing paradigms is its vision of the next generation of manufacturing with enhanced capabilities [126]. Smart manufacturing systems adapt to new circumstances by leveraging real-time information for intelligent decision-making and by proactively predicting and preventing potential failures.
@@ 1.6.6 Internet of Robotics Things (IoRT)
The Internet of Robotic Things (IoRT) is a concept that combines the principles of IoT (Internet of Things) and Robotics. IoRT represents an emerging technology that incorporates robots within an IoT ecosystem as objects, enabling communication, collaboration, and automation. These robots are seamlessly integrated into smart environments, where they perform a wide range of tasks. These tasks span from personal activities within smart homes to applications in the healthcare industry. Furthermore, they extend to professional activities such as monitoring, delivery, and object control within manufacturing industries or warehouses.
@@ 1.6.7 Oil and Gas Industry
The integration of IoT into the oil and gas industry can enable automated and optimized execution of industry operations, leading to more efficient decision-making. IoT-based applications in the oil and gas sector offer numerous advantages, including enhanced operational efficiency through automation and optimization, resulting in overall cost reduction. Environmental monitoring sensors deployed in this context can detect and alert to the presence of poisonous or harmful substances in the environment. Moreover, IoT technology provides real-time data, significantly improving decision-making processes. Importantly, IoT integration enables the early detection of potential hazards, empowering authorities to take prompt and effective preventive measures.
The oil and gas industry entails substantial expenditures and significant health and safety risks, highlighting the importance of continuous and real-time monitoring within this sector. To achieve this objective, the oil and gas industry is increasingly adopting IoT-based remote monitoring solutions. These solutions allow for the remote monitoring of field equipment, the analysis of field data, collaborative data-driven decision-making, and the implementation of control commands to optimize asset performance while mitigating health, safety, and environmental (HSE) hazards [127]. Furthermore, IoT integration in this industry aims to reduce human labor, minimize time wastage, and enhance accuracy through automation. An illustrative example is Equinor’s deployment of a well optimization system in the Bakken oil field. In this case, Equinor have utilized IoT devices and machine learning algorithms in approximately 50 wells. This implementation led to optimized well operation and maintenance, resulting in a remarkable 33% increase in oil production from these.
@@ 1.6.8 Smart Retail
The adoption of IoT in the retail industry has created a flexible environment that benefits both customers and sellers. This shift allows the entire retail sector to migrate from offline to online, enabling customers to independently conduct their shopping through self-service while facilitating smooth interactions between retailers and their customers. Furthermore, IoT technologies like RFID can be employed to monitor products, and sensors can be deployed to collect customer data, which can then be utilized for analyzing customer buying behavior and enhancing business profitability. Additionally, customers have the option to make payments through online transactions and monitor their orders using online services [128].
In the study presented in [129], the authors have introduced a smart unstaffed retail store that relies on AI and IoT for commodity detection. Furthermore, in [130], two models, namely the Retail 4.0 IoT Consumer Model and the Retail 4.0 IoT Retailer Model, have been put forward. These models are capable of analyzing consumer behavior and discerning consumer buying patterns with ease.
@@ 1.6.9 Industrial Internet of Things (IIoT)
The IIoT, often recognized as The Industrial Internet of Things (IIoT), holds significant potential according to numerous market researchers. It serves as an extension of IoT specifically customized for the industrial sector and its applications. IIoT empowers industries and enterprises to enhance and optimize their operations by leveraging machine-to-machine communication, big data analytics, and machine learning. The scope of IIoT is extensive, encompassing a wide array of connected industrial devices and systems. These include, but are not limited to, connected electric meters, wastewater systems, flow gauges, pipeline monitors, manufacturing robots, and various other types of industrial equipment and devices [131].
One notable application of IIoT is in the mining industry, where companies like CISCO have implemented IIoT solutions to improve safety and efficiency in underground mines. These solutions involve connecting people, tracking the locations of miners and vehicles, monitoring vehicle statuses, and automating building controls. Additionally, the mine’s blasting system is integrated with the personnel location tracking system to enhance miner safety [132]. Moreover, IoT has the potential to revolutionize the manufacturing field by enhancing productivity, maintenance practices, machine monitoring, and ensuring the safety of workers [133].
@@ 1.6.10 Social Life and Entertainment
In addition to work or professional activities, social life and entertainment are integral parts of a person’s life, and several applications have been developed to monitor and enhance human social activities. Portable devices such as mobile phones and tablets possess sensing capabilities and communication technologies that facilitate interactions between individuals. Integrating IoT into an individual’s social life can contribute to emotion detection, community building, and emotional support.
For instance, CircleSense [134] is an application designed to identify a person’s social circle by analyzing their social activities using various sensors. It also tracks the person’s location via location sensors and employs Bluetooth technology to identify people in proximity. Affective computing is a technology that analyzes various aspects of a person’s behavior, including facial expressions, speech, body gestures, hand movements, and sleep patterns, to identify and respond to their emotions. An application of affective computing is Camy [135]. Camy is an artificial pet dog designed to interact with human beings, displaying affection and empathy. It assesses the owner’s mood and provides suggestions, emotional support, and encourages playful behavior accordingly.
A Table 1.7 has been provided for presenting information about relevant works in the aforementioned domains.
@@
Applications Related Works
Smart home (SH) [5, 136]
Smart manufacturing system (SMS) [125, 126]
Internet of robotics things (IoRT) [137, 138]
Oil and gas [127]
Smart Retail [129, 130]
Industrial IoT [132, 133]
Social Life and Entertainment [134, 135]
@@ 1.7 Challenges and Future Directions
The remarkable position that IoT now occupies in today’s world was once only a dream a few years ago. Today, IoT has captivated the entire globe and continues to extend its reach into various domains. Nonetheless, IoT also grapples with numerous issues and challenges that pose hurdles to its seamless implementation and expansion. This section offers an in-depth exploration of the key challenges confronting IoT.
@@ 1.7.1 Broad and Open Research Challenges
Broad and open research challenges refers to complex and critical challenges in the IoT system that do no have any straightforward or predefined solutions. These challenges require extensive research, exploration in order to propose efficient solutions. They are “broad” in the sense that they encompass a wide range of related issues and considerations, and they are “open” because they may not have clear-cut or definitive answers, leaving room for ongoing research, experimentation, and discovery. This subsection provides a comprehensive listing of challenges in the IoT domain.
@@
Creating a smart environment requires a vast number of devices, sensors, and complementary technologies to facilitate their inter-connectivity. Managing this large volume of objects constitutes the initial challenge in the realm of IoT and intelligent environments. Moreover, the substantial task of collecting, storing, and conducting efficient analyses on enormous amounts of data remains a significant concern and creates collision issues within the IoT framework [139]. Handling massive amounts of devices and data in the IoT requires. One approach is to employ decentralized systems instead of centralized ones, reducing the volume of data sent to the cloud for processing. Techniques like data filtering, compression, and load balancing can further minimize data size. Utilizing IoT technologies with robust device management and maintenance capabilities is also beneficial. Additionally, leveraging big data technologies like Hadoop and Spark can efficiently handle the substantial IoT data volumes. This holistic approach ensures readiness for the expanding IoT landscape.
@@
The heterogeneity of IoT devices and the diversity of various IoT applications result in various security and privacy issues. People are primarily concerned about potential privacy invasion and security threats when using these technological devices. Further details are provided below.
@@
    Security
    Different layers of IoT are vulnerable to various kinds of attacks based on the technologies and protocols used in these respective layers. According to [140], IoT layers including perception layer, network layer, and application layer face various security attacks. As the major role of perception layer is to collect data, so the security challenges in this layer focuses in Falsifying the data and destroying perception devices. Attacks including Node Capture Attacks, Malicious code Injection Attacks, False Data Injection Attacks, Replay Attacks, Cryptanalysis Attacks & Side Channel Attacks, Eavesdropping and Interference and Sleep Deprivation Attacks are faced by this layer [140].
    Since the basic function of network layer is to transmit collected data, particularly using wireless technologies, security challenges in this layer revolve around the availability of network resources and the wireless network. Challenges at the Network Layer encompass Denial-of-Service (DoS) Attacks, Spoofing Attacks, Sinkhole Attacks, Wormhole Attacks, Man-in-the-Middle Attacks, Routing Information Attacks, Sybil Attacks, and Unauthorized Access [140]. Application layer, which focuses on providing user-requested services, challenges primarily revolve around software attacks, including: Phishing Attack, Malicious Virus/Worm, Malicious Scripts [140].
    It is imperative to propose secure, robust, and reliable authentication schemes to detect and defend unauthorized access [141]. To defend against worms, viruses, and malicious scripts bridging firewalls, virus detection techniques and script detection techniques such as honeypot techniques, static code analysis, and dynamic action detection must be implemented. Along with that, secured routing protocols must be introduced to ensure secure routing.
@@
    Privacy
    As previously mentioned, IoT devices continuously generate vast amounts of real-time data, which undergo three main stages: (1) Data Collection, (2) Data Aggregation , and (3) Data Mining and Analytics. While these processes enhance our lives by providing various services, they also raise concerns about data privacy in IoT. Privacy breaches in IoT can have serious repercussions for both the IoT network and its users, including financial losses, property damage, and even risks to human safety and security [140, 142].
    For instance, consider the smart grid, Adversaries can readily seize control of the smart meters, allowing them to access or manipulate the collected data. This could potentially compromise the confidentiality and privacy of energy consumption data. By using these altered data, utility providers may make inaccurate assessments of energy supply and demand within the grid, resulting in erroneous energy dispatch decisions. This, in turn, could lead to imbalances in energy supply and demand, potentially causing widespread power outages. In the healthcare industry, if an adversary manages to acquire a patient’s health data, they could manipulate medication prescriptions or medical records, leading to significant health risks and potential insurance fraud. Therefore, it is imperative to deploy privacy preservation schemes to prevent data leakage and ensure that private data remains inaccessible to adversaries [140, 143].
    Privacy-preserving mechanisms in the context of IoT data processing can be categorized into three main groups: (i) privacy preservation during data collection, (ii) privacy preservation during data aggregation, and (iii) privacy preservation during data mining and analytics. While various techniques such as encryption and key management can be applied to protect privacy in data collection, mining, and analytics, the majority of efforts in IoT privacy preservation have focused on data aggregation. Data aggregation involves processing relevant data in multiple locations, making it challenging to ensure privacy using traditional encryption methods. Therefore, several privacy-preserving mechanisms have been developed specifically for data aggregation, falling into the following categories:
@@
      Encryption-based privacy preservation, which utilizes encryption techniques like homomorphic encryption, commitment mechanisms, secret sharing, and zero-knowledge proofs to prevent data from being eavesdropped by adversaries during aggregation [144, 145].
@@
      Perturbation-based privacy preservation, where techniques such as data customization, data sharing, and random noise injection are used to perturb raw data, ensuring privacy during aggregation.
    Among these, perturbation-based privacy-preserving schemes are popular in IoT due to their direct operation on raw data. However, many of these perturbation-based privacy preserving schemes sacrifice data utility to achieve privacy. This reduction in data utility can hinder the support of services requested by IoT applications. Therefore, a significant challenge, in the field of data privacy preservation in IoT, is designing schemes that strike a balance between privacy and data utility, making it a crucial area for future research [21, 140]. In summary, safeguarding data privacy in IoT is crucial to prevent these adverse outcomes and maintain the security and integrity of both individuals and the IoT ecosystem.
@@
In an IoT network, devices from various vendors are interconnected, which can lead to monitoring and management challenges. Different industries currently rely on a multitude of standards to support their applications. Given the vast amounts of data, diverse device types, and the presence of various entities, utilising standard interfaces becomes crucial. This importance is amplified, especially for applications that need to accommodate both cross-organisational collaborations and a wide array of system limitations. Addressing these issues requires all industries to adhere to specific standards, but achieving such universal compliance can be a daunting and impractical task.
@@
IoT is an ever-expanding network, and heterogeneous devices are expected to continuously join the IoT network in the future. As a result, as the number of devices increases, ensuring smooth connectivity, effective data management, and overall system performance at a small scale becomes increasingly challenging. Therefore, the scalability of IoT poses an ongoing challenge for the future of this technology. To effectively address scalability challenges, it is essential to construct a scalable architecture, utilizing technologies like modular components, load balancers, and distributed systems.
@@
Small smart devices that make up IoT systems often have limited battery power, which cannot be easily replaced. This limitation can lead to a global energy crisis and high power consumption, as well as constraints on memory and processing capabilities. Consequently, routing processes and compute-intensive applications may not run efficiently on these devices. While some routing protocols do support low-power communication, they are still in the early stages of development, and the constrained energy of smart devices may not be sufficient to fully utilize these WSN routing protocols. To tackle these challenges effectively, it’s essential to emphasize the creation of low-power hardware and the adoption of energy-efficient protocols like MQTT-SN or CoAP, rather than relying on more power-intensive alternatives such as HTTP. Additionally, harnessing over-the-air (OTA) firmware updates can ensure devices remain optimized and bug-free, thus diminishing the necessity for physical maintenance visits. Duty cycling, as another viable approach, aids in curbing power consumption significantly.
@@
Mobility management in IoT refers to the ability to handle devices that move within the network seamlessly. It is a crucial aspect because many IoT devices are not stationary and need to communicate as they change locations. The presence of mobile devices in IoT setups can lead to challenges in how routing protocols and IoT networks work efficiently. The current methods used for devices that move, like in sensor networks, mobile adhoc networks, and vehicular networks, can’t effectively handle the various problems related to routing because of these sensors having limited processing power and energy resources. To address these challenges, IoT systems employ various mobility management techniques and protocols, aiming to provide reliable and seamless communication for mobile devices in the IoT ecosystem.
@@
The procedure for Processing, Analysis and Management of Data is tremendously challenging because of the heterogeneous nature of IoT devices and large sale of data generation. Currently, most systems utilize the centralized cloud based system for performing computationally intensive tasks and delivery of data. However, an ongoing concern revolves around the limitations of traditional cloud architectures when it comes to efficiently handling the vast amounts of data generated and utilized by IoT-enabled devices. Additionally, these architectures struggle to support the associated computational demands while also meeting precise timing constraints. To address this challenge, most systems are currently relying on existing solutions like mobile cloud computing and fog computing, both of which utilize edge processing [146].

### 12 Jan 2024 v2

@@ 1.6.4 Smart Home (SH)
In a smart home, various types of sensors are strategically deployed, each with its own specific function. Smart homes simplify daily tasks for users, proving especially beneficial for those prone to forgetting routine actions like locking doors or turning off appliances. From smart door locks to the maintenance of household items like coffee machines, heaters, and smart bulbs, and even the use of surveillance cameras for enhanced security, smart homes offer a wide range of possibilities. Furthermore, users can control these devices through voice commands and remotely monitor their home equipment. Smart homes contribute to improved energy efficiency by automatically turning off devices not in use and notifying users of any unusual incidents. MavHome [5], for example, employs prediction algorithms to perform various tasks in response to user-initiated events. As for energy conservation, an intelligent home achieves it through the utilization of sensors and the context-aware capabilities of IoT. Data gathered by these sensors is transmitted to a context aggregator, which then forwards the data to a context-aware service engine. This engine analyzes the data and determines appropriate actions. For instance, it may decide to turn off the air conditioning if the temperature is too cold, shut off the gas supply in case of a detected leak, or switch off the lights when there are no occupants at home [124].
@@ 1.6.5 Smart Manufacturing System (SMS)
With the development and evolution of IoT, industrial IoT, artificial intelligence, and cyber-physical systems, many countries have opted to transform their manufacturing systems into smart manufacturing systems. Through the integration of smart technologies, these systems facilitate a rapid and extensive flow of data within and among manufacturing processes. Equipped with this data and employing advanced information and communication technology, smart manufacturing systems possess the capability to swiftly respond to global demands, efficiently utilise materials, energy, and labour resources, and deliver customised products on time [125]. What sets the smart manufacturing model apart from other manufacturing paradigms is its vision of the next generation of manufacturing with enhanced capabilities [126]. These systems adapt to new circumstances by leveraging real-time information for intelligent decision-making and by proactively predicting and preventing potential failures.
@@ 1.6.6 Internet of Robotics Things (IoRT)
The Internet of Robotic Things is a concept that combines the principles of IoT and robotics. IoRT represents an emerging technology that incorporates robots within an IoT ecosystem as objects, enabling communication, collaboration, and automation. These robots seamlessly integrate into smart environments, performing a wide range of tasks. These tasks span from personal activities within smart homes to applications in the healthcare industry. Furthermore, they extend to professional activities such as monitoring, delivery, and object control within manufacturing industries or warehouses.
@@ 1.6.7 Oil and Gas Industry
The oil and gas industry, facing substantial costs and safety risks, embraces IoT-based remote monitoring for real-time field equipment oversight and data-driven decision-making. These solutions allow for the remote monitoring of field equipment, the analysis of field data, collaborative data-driven decision-making, and the implementation of control commands to optimize asset performance while mitigating health, safety, and environmental (HSE) hazards [127]. Furthermore, IoT integration in the oil industry focuses on reducing human labor, minimizing time wastage, and improving accuracy through automation, as exemplified by Equinor’s well optimization system in the Bakken oil field [freeman2018enabling]. By deploying IoT devices and ML algorithms in around 50 wells, Equinor achieved a 33% increase in oil production through optimized well operation and maintenance.
@@ 1.6.8 Smart Retail
The adoption of IoT in the retail industry has created a flexible environment that benefits both customers and sellers. This shift allows the entire retail sector to migrate from offline to online, enabling customers to independently conduct their shopping through self-service while facilitating smooth interactions between retailers and their customers. Furthermore, retailers can employ IoT technologies like RFID to monitor products and deploy sensors to collect customer data, which they can then utilize for analyzing customer buying behavior and enhancing business profitability [129, 130]. Additionally, customers have the option to make payments through online transactions and monitor their orders using online services [128].
@@ 1.6.9 Industrial Internet of Things (IIoT)
The IIoT holds significant potential, according to numerous market researchers, by serving as an extension of the IoT specifically customized for the industrial sector and its applications. It empowers industries and enterprises to enhance and optimize their operations by leveraging M2M communication, big data analytics, and ML. The scope of IIoT is extensive, encompassing a wide array of connected industrial devices and systems. Connected electric meters, wastewater systems, flow gauges, pipeline monitors, manufacturing robots, and various other types of industrial equipment and devices are included in this list [131]. One notable application of IIoT is in the mining industry, where companies like CISCO have implemented IIoT solutions to improve safety and efficiency in underground mines. These solutions involve connecting people, tracking the locations of miners and vehicles, monitoring vehicle statuses, and automating building controls.
@@ 1.6.10 Social Life and Entertainment
Several applications have been developed to monitor and enhance human social activities, in addition to work or professional activities, as social life and entertainment are integral parts of a person’s life. Portable devices such as mobile phones and tablets possess sensing capabilities and communication technologies that facilitate interactions between individuals. Integrating IoT into an individual’s social life can contribute to emotion detection, community building, and emotional support. CircleSense [134] is an application that analyzes a person’s social activities using various sensors to identify their social circle. It also tracks the person’s location via location sensors and employs Bluetooth technology to identify people in proximity. Camy, an artificial pet dog, expresses affection and empathy through the use of effective computing technology. This technology analyzes multiple aspects of a person’s behavior, such as facial expressions, speech, body gestures, hand movements, and sleep patterns, to identify and appropriately respond to their emotions [135]. A Table 1.8 has been provided for presenting information about relevant works in the aforementioned domains.
@@
Applications Related Works Applications Related Works
Social Life and
Entertainment
[134, 135]
Smart manufacturing
system
[125, 126]
Internet of robotics things [137, 138] Oil and gas [127]
Smart Retail [129, 130] Industrial IoT [132, 133]
Smart home [5, 136]
@@ 1.7 Challenges and Future Directions
The remarkable position that IoT now occupies in today’s world was once only a dream a few years ago. Today, IoT has captivated the entire globe and continues to extend its reach into various domains. Nonetheless, IoT also grapples with numerous issues and challenges that pose hurdles to its seamless implementation and expansion. This section offers an in-depth exploration of the key challenges confronting the IoT.
@@ 1.7.1 Broad and Open Research Challenges
Broad and open research challenges refer to complex and critical challenges in the IoT system that do not have any straightforward or predefined solutions. These challenges require extensive research and exploration in order to propose efficient solutions. They are “broad” in the sense that they encompass a wide range of related issues and considerations, and they are “open” because they may not have clear-cut or definitive answers, leaving room for ongoing research, experimentation, and discovery. This subsection provides a comprehensive listing of challenges in the IoT domain.
@@
Creating a smart environment requires a vast number of devices, sensors, and complementary technologies to facilitate their interconnectivity. Managing this large volume of objects constitutes the initial challenge in the realm of IoT and intelligent environments. Moreover, the substantial task of collecting, storing, and conducting efficient analyses on enormous amounts of data remains a significant concern and creates collision issues within the IoT framework [139]. Handling massive amounts of devices and data in the IoT requires. One approach is to employ decentralized systems instead of centralized ones, reducing the volume of data sent to the cloud for processing. Techniques like data filtering, compression, and load balancing can further minimize the size of the data. Utilizing IoT technologies with robust device management and maintenance capabilities is also beneficial. Additionally, leveraging big data technologies like Hadoop and Spark can efficiently handle the substantial IoT data volumes. This holistic approach ensures readiness for the expanding IoT landscape.
@@
The heterogeneity of IoT devices and the diversity of various IoT applications result in various security and privacy issues. People are primarily concerned about potential privacy invasions and security threats when using these technological devices. More information is provided below.
@@
    Security: Different layers of the IoT are vulnerable to various kinds of attacks based on the technologies and protocols used in these layers. According to [140], IoT layers including perception layer, network layer, and application layer face various security attacks. As the major role of the perception layer is to collect data, the security challenges in this layer focus on falsifying the data and destroying perception devices. Attacks including Node Capture Attacks, Malicious code Injection Attacks, False Data Injection Attacks, Replay Attacks, Cryptanalysis Attacks and Side Channel Attacks, Eavesdropping and Interference, and Sleep Deprivation Attacks are faced by this layer [140].
    Since the basic function of network layer is to transmit collected data, particularly using wireless technologies, security challenges in this layer revolve around the availability of network resources and the wireless network. Challenges at the Network Layer encompass Denial-of-Service (DoS) Attacks, Spoofing Attacks, Sinkhole Attacks, Wormhole Attacks, Man-in-the-Middle Attacks, Routing Information Attacks, Sybil Attacks, and Unauthorized Access [140]. Application layer, which focuses on providing user-requested services, challenges primarily revolve around software attacks, including phishing attacks, malicious viruses/worms, malicious scripts [140].
    It is imperative to propose secure, robust, and reliable authentication schemes to detect and defend unauthorized access [141]. Virus detection techniques and script detection techniques such as honeypot techniques, static code analysis, and dynamic action detection must be implemented to defend against worms, viruses, and malicious scripts bridging firewalls. Furthermore, introducing secured routing protocols is essential to ensure secure routing.
@@
    Privacy: IoT devices continuously generate vast amounts of real-time data, which undergoes three main stages: (1) Data Collection, (2) Data Aggregation , and (3) Data Mining and Analytics. While these processes enhance our lives by providing various services, they also raise concerns about data privacy in the IoT. Privacy breaches in IoT can have serious repercussions for both the IoT network and its users, including financial losses, property damage, and even risks to human safety and security [140, 142].
    For instance, consider the smart grid, where adversaries can readily seize control of the smart metres, allowing them to access or manipulate the collected data. This could potentially compromise the confidentiality and privacy of energy consumption data. By using this altered data, utility providers may make inaccurate assessments of energy supply and demand within the grid, resulting in erroneous energy dispatch decisions. This, in turn, could lead to imbalances in energy supply and demand, potentially causing widespread power outages. In the healthcare industry, if an adversary manages to acquire a patient’s health data, they could manipulate medication prescriptions or medical records, leading to significant health risks and potential insurance fraud. Therefore, it is imperative to deploy privacy preservation schemes to prevent data leakage and ensure that private data remains inaccessible to adversaries [140, 143].
    There are three main groups for categorizing privacy-preserving mechanisms in the context of IoT data processing: (i) privacy preservation during data collection, (ii) privacy preservation during data aggregation, and (iii) privacy preservation during data mining and analytics. While various techniques, such as encryption and key management, can be applied to protect privacy in data collection, mining, and analytics, the majority of efforts in IoT privacy preservation have focused on data aggregation. Data aggregation involves processing relevant data in multiple locations, making it challenging to ensure privacy using traditional encryption methods. As a result, researchers have developed several privacy-preserving mechanisms specifically for data aggregation, which can be categorized as follows:
@@
      Encryption-based privacy preservation, which prevents adversaries from eavesdropping on data during aggregation by utilizing encryption techniques such as homomorphic encryption, commitment mechanisms, secret sharing, and zero-knowledge proofs [144, 145].
@@
      Perturbation-based privacy preservation, where techniques such as data customization, data sharing, and random noise injection perturb raw data to ensure privacy during aggregation.
    Among these, perturbation-based privacy-preserving schemes are popular in IoT due to their direct operation on raw data. However, many of these perturbation-based privacy-preserving schemes sacrifice data utility to achieve privacy. This reduction in data utility can hinder the support of services requested by IoT applications. Therefore, a significant challenge in the field of data privacy preservation in the IoT is designing schemes that strike a balance between privacy and data utility, making it a crucial area for future research [21, 140]. In summary, safeguarding data privacy in the IoT is crucial to preventing these adverse outcomes and maintaining the security and integrity of both individuals and the IoT ecosystem.
@@
Interconnecting devices from various vendors in an IoT network can pose monitoring and management challenges. Different industries currently rely on a multitude of standards to support their applications. Given the vast amounts of data, diverse device types, and the presence of various entities, utilising standard interfaces becomes crucial. This importance is amplified, especially for applications that need to accommodate both cross-organisational collaborations and a wide array of system limitations. Addressing these issues requires all industries to adhere to specific standards, but achieving such universal compliance can be a daunting and impractical task.
@@
In the future, heterogeneous devices are expected to continuously join the ever-expanding IoT network. As a result, as the number of devices increases, ensuring smooth connectivity, effective data management, and overall system performance on a small scale becomes increasingly challenging. Therefore, the scalability of IoT poses an ongoing challenge for the future of this technology. To effectively address scalability challenges, it is essential to construct a scalable architecture, utilizing technologies like modular components, load balancers, and distributed systems.
@@
Small smart devices that comprise IoT systems often have limited battery power, which is not easily replaceable. This limitation can lead to a global energy crisis and high power consumption, as well as constraints on memory and processing capabilities. Consequently, routing processes and compute-intensive applications may not run efficiently on these devices. While some routing protocols do support low-power communication, they are still in the early stages of development, and the constrained energy of smart devices may not be sufficient to fully utilize these WSN routing protocols. To tackle these challenges effectively, it’s essential to emphasize the creation of low-power hardware and the adoption of energy-efficient protocols like MQTT-SN or CoAP rather than relying on more power-intensive alternatives such as HTTP. Additionally, harnessing over-the-air (OTA) firmware updates can ensure devices remain optimized and bug-free, thus diminishing the necessity for physical maintenance visits. Duty cycling, as another viable approach, aids in curbing power consumption significantly.
@@
Mobility management in the IoT refers to the ability to handle devices that move within the network seamlessly. It is a crucial aspect because many IoT devices are not stationary and need to communicate as they change locations. The presence of mobile devices in IoT setups can lead to challenges in how routing protocols and IoT networks work efficiently. The current methods used for devices that move, like in sensor networks, mobile adhoc networks, and vehicular networks, can’t effectively handle the various problems related to routing because these sensors have limited processing power and energy resources. To address these challenges, IoT systems employ various mobility management techniques and protocols, aiming to provide reliable and seamless communication for mobile devices in the IoT ecosystem.
@@
The procedure for processing, analysing, and managing data is tremendously challenging because of the heterogeneous nature of IoT devices and the large scale of data generation. Currently, most systems utilize the centralized cloud-based systems for performing computationally intensive tasks and delivering data. However, an ongoing concern revolves around the limitations of traditional cloud architectures when it comes to efficiently handling the vast amounts of data generated and utilized by IoT-enabled devices. Additionally, these architectures struggle to support the associated computational demands while also meeting precise timing constraints. To address this challenge, most systems are currently relying on existing solutions like mobile cloud computing and fog computing, both of which utilize edge processing [146].
Other Challenges
In addition to the previously mentioned challenges, IoT technology faces several other issues. The widespread adoption of IoT devices and technology, coupled with our increasingly reliant lifestyles, has led to users becoming highly dependent on IoT applications. This reliance is particularly critical in healthcare, where patients heavily depend on healthcare applications. Moreover, IoT devices can sometimes unexpectedly interfere with human activities, resulting in unanticipated and autonomous behaviors. The IoT network introduces ambiguity, making it challenging to distinguish between physical and virtual devices and even humans due to the ease of transformation between these categories. Quality and traffic control have become more complex due to the miniaturization and huge number of IoT devices. Managing unique identifications for each IoT device is also a growing concern. Furthermore, the IoT goes beyond geographical boundaries, with applications like healthcare offering services internationally. Nations face challenges due to the global reach of IoT, as data generated within their borders can be collected and transmitted to service providers located anywhere in the world, giving rise to concerns regarding data privacy and jurisdiction. Addressing these multifaceted challenges will require careful consideration and international collaboration to ensure the effective and secure implementation of IoT technology.
1.7.2 Ethical Considerations
The term “ethical issue” pertains to a situation or quandary characterized by a clash of moral principles, values, or ethical norms [147]. The realm of IoT confronts these ethical quandaries, necessitating individuals or organizations to make challenging choices amidst conflicting interests. These decisions frequently revolve around determining what is ethically correct or incorrect. Figure 1.6 illustrates prevalent ethical dilemmas within the IoT domain.